Skysong Inc. issues 500 shares of $10 par value common stock and 100 shares of $100 par value preferred stock for a lump sum of $123,000.

(a) Prepare the journal entry for the issuance when the market price of the common shares is $168 each and market price of the preferred is $210 each.

(b) Prepare the journal entry for the issuance when only the market price of the common stock is known and it is $216 per share.

A)

Acc DR CR

B)

ACC Dr Cr
Answer #1

Market value of Common Shares = 168*500 = 84000 : Market value of preferred shares = 100*210 = 21000

a) Journal entry

Acc Dr Cr
Cash 123000
Common Stock 5000
Paid in capital in excess of par-Common Stock (123000*84/105)-5000 93400
Preferred stock (100*100) 10000
Paid in capital in excess of par-Preferred stock (123000*21/105)-10000 14600

b) Journal entry

Acc Dr Cr
Cash 123000
Common Stock 5000
Paid in capital in excess of par-Common Stock (500*216)-5000 103000
Preferred stock (100*100) 10000
Paid in capital in excess of par-Preferred stock 5000